# Beyond correlation: The impact of human uncertainty in measuring the effectiveness of automatic evaluation and LLM-as-a-judge

### v2

@@ Abstract
The effectiveness of automatic evaluation of generative models is typically measured by comparing it to human evaluation using correlation metrics. However, metrics like Krippendorff’s $\alpha$ and Randolph’s $\kappa$, originally designed to measure the reliability of human labeling, make assumptions about human behavior and the labeling process. In this paper, we show how relying on a single aggregate correlation score can obscure fundamental differences between human behavior and automatic evaluation methods, including LLM-as-a-Judge. Specifically, we demonstrate that when the proportion of samples with variation or uncertainty in human labels (gathered during human evaluation) is relatively high, machine labels (generated by automatic evaluation methods) may superficially appear to have similar or better correlation with the human majority label compared to human-to-human (HH) correlation. This can create the illusion that automatic evaluation approximates the human majority label. However, as the proportion of samples with consistent human labels increases, the correlation between machine and human labels fall well below HH correlation. Based on these findings, we first propose stratifying results by human label uncertainty to provide a more robust analysis of automatic evaluation performance. Second, recognizing that uncertainty and variation are inherent in perception-based human evaluations, such as those involving attitudes or preferences, we introduce a new metric - binned Jensen-Shannon Divergence for perception for such scenarios to better measure the effectiveness of automatic evaluations. Third, we present visualization techniques – perception charts, to compare the strengths and limitations of automatic evaluation and to contextualize correlation measures appropriately.
@@ 1 Introduction
Despite the increasing importance of automatic evaluation for generative large language models (LLMs), measuring and interpreting the effectiveness of these methods remains challenging. Automatic evaluation methods include $N$-gram-based methods such as Rouge (Lin, 2004) and BERTScore (Zhang* et al., 2020), task-specific model-based methods (TM) such as AlignScore (Zha et al., 2023) and Prometheus (Kim et al., 2024), and LLM-as-a-Judge-based methods (LJ) where a general purpose LLM, using prompts, is instructed to perform evaluation. The use of automatic evaluation is tied to the key question – How can the efficacy of automatic evaluation methods be measured in comparison to human evaluation? A widely accepted practice is to use agreement or rank correlation metrics such as Krippendorff’s-$\alpha$ (Krippendorff, 2011), Cohen’s-$\kappa$ (Cohen, 1960; McHugh, 2012), Spearman’s $\rho$, and Kendall’s $\tau$-b (Fabbri et al., 2021; Deutsch et al., 2022; Chiang & Lee, 2023; Liu et al., 2023b), where a higher correlation between machine labels (generated by automatic evaluation methods) and human labels (gathered during human evaluation) implies a more effective automatic evaluation. Variation or uncertainty in human labels (Tversky & Kahneman, 1974) is almost impossible to avoid when humans are involved (Kahneman et al., 2021). In this paper, we first highlight how comparing the overall correlation score between human labels and machine labels can obscure key discrepancies in machine generated labels. Therefore, we propose that measurements be stratified by label uncertainty.
@@
Secondly, this problem also leads to a follow-on question about how to interpret uncertainty. Not all types of uncertainties are equal (Kahneman & Tversky, 1982), as they can range from random human errors, ambiguity in tasks, and nature of perception-based evaluation (Elangovan et al., 2024), where some of these variations are inherent to the task while others indicate poor reliability of human labels. For instance, when evaluating aspects such as “How readable is the given content?”, the answer relies on human perception and is likely to vary from person to person. In contrast, evaluating a natural language inference (NLI) task, such as “Can you infer the given sentence from a given text?”, is less likely to show variations. Acceptable variation in labels, therefore, depends on the task, as variations or uncertainties in human labels can sometimes be a feature rather than a bug in human evaluation. Thus, we explore how to effectively measure and interpret reliability in the context of human uncertainties. Hence, we ask the following questions in this paper
@@
  • Research Question 1 - How does uncertainty in human labels impact correlation metrics when we measure the efficacy of automatic evaluation methods? (Sec. 3.1)
@@
  • Research Question 2 — How can we measure human-to-machine (HM) agreement that accounts for human uncertainty as a result of variation in human perception? (Sec. 3.2)
@@
  • Research Question 3 — How can we visualize the underlying data to draw meaningful insights when we compare the results from automatic and human evaluations? (Sec. 3.3)
@@ 2.1 Correlation measurements and assumptions
Correlation measures have traditionally been used to measure the inter-rater agreement (IRA) of human-generated labels, where there is no ground truth available, to measure human-to-human (HH) correlation. There are 2 broad categories of correlation measures, (a) non-chance-adjusted indices and (b) chance-adjusted metrics (Xinshu Zhao & Deng, 2013). Chance-adjusted measures take into account a scenario where raters can choose the same label by chance even when they do not know the correct answer. Chance-adjusted measures, such as Krippendorff’s-$\alpha$ are generally more popular IRA measures than non-chance-adjusted metrics such as percentage agreement. The key difference between various chance-adjusted IRA measures is how chance agreement is computed (Meyer et al., 2014; Xinshu Zhao & Deng, 2013). For instance, Randolph’s-$\kappa$ (Randolph, 2005) assumes that all labels are equally likely and assigns a probability of $\nicefrac{{1}}{{k}}$ for chance agreement, where $k$ is the number of possible label choices. On the other hand, Fleiss’s $\kappa$ (Fleiss, 1971) and Krippendorff’s $\alpha$ calculate chance agreement empirically, based on the actual human labels of the study. These measurements can be useful in cases where certain labels (e.g., 3 on a Likert scale of 1-5) are more likely to be chosen when humans are uncertain. A high proportion of such labels, even if consistent, may indicate unreliability in the labeling process, particularly when the minority labels show lower agreement. Therefore, chance-adjusted correlation measures rely on key assumptions about the types of errors humans are likely to make. Furthermore, Krippendorff’s-$\alpha$, specifies that the agreement coefficient is an indicator of reliability if and only if (1) the raters duplicate the process of rating, i.e., when two independent raters use the same guidelines or instructions and use the same units, (2) raters must be treated as interchangeable, and (3) correct answers are not known (Krippendorff, 2004). Human labels and machine labels are clearly not interchangeable even when using current state-of-the-art LLMs to generate machine labels, as humans and LLMs are fundamentally different.
@@ 2.2 Differences between correlation measures and raw accuracy
The assumptions above, especially the assumption that raters must be treated as interchangeable, raise a fundamental question as to why even use correlation measures to compare machine and human performance, instead of measures such as accuracy and F1 using human labels as ground truth? Firstly, one of the main challenges with using scores like accuracy or F1 is that human labels tend to be noisy regardless of the task or the dataset, where some label choices are more acceptable than others, and measures such as F1 or accuracy do not account for variations in human labels and hence correlation measures appear to be a better measure to compare human vs. machine performance. However, as mentioned previously, many of the common chance-adjusted correlation measures aim to account for random chance agreement in HH agreement, whereas the errors make by machines tend to be more systematic. A key characteristic of random errors is that they are non-reproducible which result in variations when repeated, resulting in low IRA, which in-turn may point to potential problems in the labelling process or complexities in the task. In contrast, systematic errors are consistent, and IRA measures are not designed to detect them. Secondly, as a direct consequence, IRA metrics assume that agreement is good and variance is bad, which implies that there is a single correct answer – a situation that may not hold for ambiguous items or those that gauge user perception.
@@ 2.3 Types of raw data collected during evaluation
The type of the data collected, such as nominal, ordinal, interval, ratio, discrete or continuous data, dictate how such data can be aggregated and interpreted. Nominal and ordinal measures can be considered as qualitative measurements, while continuous values can be interpreted as quantitative (Mishra et al., 2018). The most common types of values collected during automatic or human evaluation are:
@@
Ordinal values: Ordinal values have an implicit order, e.g., ratings of quality (very good, good, fair, poor, very poor) (Mishra et al., 2018). The most common use of ordinal values in human evaluation are through Likert scales. Likert scales are rating systems of attitude scale to measure the degree to which a person agrees/disagrees with a given statement (Taherdoost, 2019) and commonly to gauge user perception on qualitative aspects such as readability, coherence, and fluency (van der Lee et al., 2021). IRA measures for ordinal values should ideally account for the fact that ratings that are closer to one another (1 vs. 2) for a given item indicate a higher agreement than ratings that are further apart (1 vs. 4). Such IRA measures include Krippendorff’s-$\alpha$ for ordinal values, where disagreements are weighted by distance between the ratings. Rank correlations such as Kendall’s-$\tau$ and Spearman’s-$\rho$ are also widely used to measure item-level correlation between machines and humans for ordinal values. However, there’s a caveat on using these measures: their ability to handle a relatively large proportion of ties is widely debated, even when measuring system correlations that involve a much smaller proportion of ties (Deutsch et al., 2023). When the number of items to rank is $n$, the number of ordinal values to assign is $r$, and when $r<<n$, it results in many ties or rank collision. For instance, when assigning a Likert scale of 1-5 to 100 items, at least 20% of the items will have the same rank. While ordinal values preserve order, the objective of using ordinal values is not to rank the item in relation to other items, therefore the proportion of ties affects the interpretability and reliability of the rank correlation scores.
@@ 3 Analysis Settings and Results
Datasets: We use datasets that have multiple human annotations used in typical evaluation tasks for (a) Likert-based ordinal qualitative evaluation, (b) NLI datasets for fact or consistency checking, and (c) preference datasets for understanding human preferences as follows: 1. SummEval (Fabbri et al., 2021) is human evaluation of summary quality using four criteria – fluency, coherence, consistency, and relevance, with Likert scores assigned from 1 to 5. This dataset has 1600 samples, where each item is labeled by 3 expert annotators. 2. SNLI (Bowman et al., 2015) is an NLI inference task dataset with 10K samples and 3 labels – entailment, neutral and contradiction. Each item is assigned 5 labels by human annotators. 3. MNLI (Williams et al., 2018) is also an NLI inference task dataset with 10K samples and 3 labels – entailment, neutral and contradiction. Each item is assigned 5 labels by human annotators. It contains two splits – matched and mismatched. 4. MT-Bench (Zheng et al., 2023) is a preference dataset which compares human preferences of several model outputs on open-ended questions using multi-turn conversations. From this dataset, we only use samples that have at least 3 human ratings.
Models: For SummEval dataset, we reuse the original G-Eval results (Liu et al., 2023b) which rates the quality of summaries on a scale of 1-5. In addition, we also use Claude Sonnet 3 and Mistral Large v2 on SummEval, re-using the prompts from G-Eval as detailed in Appendix A.2. For NLI and MNLI datasets, we use Llama 3.1 8B, Mistral Large v2 and Sonnet 3, reusing the prompts from Liu et al. (2023a) as detailed in Appendix A.1. For MTBench, we rely on the existing results on preference data on MT-Bench and GPT-4 from Zheng et al. (2023). We also adapt the same prompts (see Appendix A.3) for Sonnet as LJ for MTBench. In order to compute machine-to-machine (MM) correlation, we set the sampling temperature to 1.0 to maximize the variation across runs, and set top p = 0.8. We sample 20 annotations per example. NOTE: We have predominately reused the prompts from existing works primarily relying on GPT-4 as LJ, we have NOT attempted to optimize the prompts per LLM except for minor format changes. Since models are known to be highly sensitive to prompts, our experiments in this paper are not meant to rank the underlying LJ, but to study the impact of human uncertainty regardless of the LJ.
@@
To study the impact of human uncertainty, we stratify samples and then compare HH and HM correlation measurements for each group. We primarily stratify by percentage agreement, which is the proportion of labels that reflect the majority (for nominal data) or the median label (for ordinal data) among all human labels assigned to an item. For example, if 3 out of 5 people assign the same label, the percentage agreement is 60%. We also stratify samples by the number of unique human labels to ensure that our findings are consistent regardless of the stratification method.
At surface level, HM correlation seems to improve with human certainty, as shown in Table 3.1 – column $H^{w}M^{w}$ comparing Human majority ($H^{w}$) with machine majority ($M^{w}$). On closer inspection, another trend emerges where in any given stratified group when the noisy or uncertain samples increases (as measured by low HH correlation), the HM correlation seems to outperform HH correlation as highlighted. However, as the proportion of samples with noisy labels decreases and HH correlation approaches near-perfect agreement, HH correlation is much higher than HM as highlighted . Thus, the illusion that machines (specifically LJs) approximate majority human behavior is largely due to the presence of noisy labels in the data. As a corollary, if LJs were truly approximating the human majority, the $\Delta$ would approach 0 under perfect HH correlation – which is clearly not the case. We also replicated this behavior on SummEval dataset which relies on ordinal labels in Table 2, using correlation metrics such as Krippendorff’s-$\alpha$ for ordinal data, Kendall’s-$\tau$ and Spearman’s-$\rho$. The same patterns are observed on preference data in Table 3.
@@
Table 1: Compare HM and HH correlation measurements on NLI datasets using Sonnet to generate machine labels. The overall correlation scores (All) indicate that the HM correlation is similar to HH. However, when stratifying by percentage agreement (PA) or human unique labels, the difference ($\Delta=HH-\text{H}^{w}\text{M}^{w}$) is highest when the all human labels are the same (PA=1), and $\text{H}^{w}\text{M}^{w}$ outperforms HH when certainty is the lowest. $H^{w}$ and $M^{w}$ are human and machine majority label.
@@
Figure 1: Krippendorff’s $\alpha$ (K-$\alpha$) on subsets of MNLI-mismatched with varying uncertainty, using 3 LLMs – Sonnet, Mistral, and Llama. As proportion of samples with consistent labels increases (towards the right end of X-axis), HH correlations are significantly higher than $\text{H}^{w}\text{M}^{w}$. $\text{H}^{w}\text{H}^{\prime}$ represents the correlation between the human majority and another human annotator ($\text{H}^{\prime}$). $\text{H}^{w}\text{R}^{w}$ compares the human majority with a random machine evaluator, serving as baseline performance.
Figure 2: Simulating the impact of uncertainty by comparing with an automatic random labeler (R). When the proportion of samples with uncertainty is higher, even a random labeler can appear to have better correlation with a majority ($\text{H}^{w}$) or median $\overline{H}$ human label. As the proportion of samples with consistency increases, the weakness of the random labeler become evident.
We further randomly sample subsets from NLI datasets (sampling detailed in Appendix 5) and compare correlation scores across subsets with varying levels of human label uncertainty. We repeat this for three different LLMs, and consistently observe the same findings, performance gaps become more apparent as human label consistency increases, shown in Fig. 1. Similar findings for the SummEval dataset are shown in Appendix Fig. 5.
Exploring the phenomena using a random labeler in a synthetically labeled dataset (code in A.10), that simulates a binary classification as well as an ordinal label {1,2,3} classification dataset, where the simulator (mimicking 2 humans) assigns 2 labels to each item. The random labeler simply assigns a random label. We plot the correlation between the random labeler and simulated humans across varying proportions of certainty. In this scenario, even a random labeler can appear better when the proportion of samples with uncertainty is higher. It is only when the proportion of samples with consistent labels increase, does the relative poor performance of the random labeler come to light, shown in Fig. 2. Intuitively, if 2 humans disagree, a 3rd random labeler cannot do any worse in-terms of agreement. The random labeler can only disagree (in which case they are no better than the 2 humans) or agree with any chosen label. This example further illustrates why stratification by proportion of uncertainty is crucial to uncovering weaknesses in any automatic labeler.
@@
Table 2: Results on SummEval dataset for criteria – Coherence and Relevance, $\overline{H}$ is the human median rating and $\overline{M}$ is machine median. $\Delta=(HH-\overline{H}\overline{M}$).Impact of human uncertainty: When the human label uncertainty is high , $\overline{H}\overline{M}$ can seem better than HH correlation and vice versa under low uncertainty . Does the correlation score ($\rho$ of 0.56 , $\alpha$ 0.49) imply the LJ is good enough to replace human evaluation? Spearman’s-$\rho$ median $(\overline{H}\overline{M}$) vs. mean ($H^{\mu}M^{\mu}$) : Median vs. mean results in an 11 points boost 0.49 $\rightarrow$ 0.60. Differences in correlation metrics: Krippendorff’s-$\alpha$ of 0.25 (reject results), while Spearman’s-$\rho$ shows 0.53 (moderate correlation).LJ ranking and absolute correlation numbers change depending on reference human under high uncertainty, Mistral best performance 0.42 replaced by Sonnet 0.22 when reference human is $H^{\prime}$. Full table in Appendix 5.
@@
Table 3: Stratified performance on preference data with at least 3 human ratings from MT-Bench for model pairs (MP). Impact of human label uncertainty: When the HH label uncertainty is high, $\text{H}^{w}\text{M}^{w}$ can seem better than HH correlation and vice versa. IRA vs. Win-rate: While the difference in win-rate between model and human ($\Delta$) can be smaller in one evaluator model (e.g., Sonnet 0.08) compared to another (e.g., GPT 0.21), the $\text{H}^{w}\text{M}^{w}$ IRA can tell an opposite story, as in the example above, GPT has better correlation with human majority across all correlation metrics except $JS_{b}$. LJ ranking and absolute correlation numbers change depending on reference human $H^{\prime}$, mimicking single annotator, under high uncertainty compared to using consensus human majority ($H^{w}$).
3.2 RQ2: How can we measure human-machine agreement that account for human uncertainty as a result of variation in human perception?
@@
where for ordinal values, bin $\mathbf{B}_{r}\coloneqq\{i\in\{1,..n\}|\text{median}(H_{i})=r\}$ contains the set of samples with a median human rating of $r$. For nominal perception-based preference labels, bin $\mathbf{B}_{r}\coloneqq\{i\in\{1,..n\}|\text{mostfrequent}(H_{i})=r\}$ contains the set of samples with the most frequent human rating of $r$.

### v3

@@ Abstract
The effectiveness of automatic evaluation of generative models is typically measured by comparing the labels generated via automation with labels by humans using correlation metrics. However, metrics like Krippendorff’s $\alpha$ and Randolph’s $\kappa$ were originally designed to measure the reliability of human labeling, thus make assumptions about typical human labeling behavior, and these assumptions may not be applicable to machine generated labels. In this paper, we show how relying on a single aggregate correlation score can obscure fundamental differences between human labels and those from automatic evaluation, including LLM-as-a-Judge. Specifically, we demonstrate that when the proportion of samples with variation or uncertainty in human assigned labels is relatively high, machine labels (generated by automatic evaluation methods) may superficially appear to have similar or better correlation with the human majority label compared to the human-to-human (HH) correlation. This can create the illusion that labels from automatic evaluation approximates the human majority label. However, as the proportion of samples with consistent human labels increases, the correlation between machine and human labels fall well below HH correlation. Based on these findings, we first propose stratifying data by human label uncertainty to provide a more robust analysis of automatic evaluation performance. Second, recognizing that uncertainty and variation are inherent in perception-based human evaluations, such as those involving attitudes or preferences, we introduce a new metric - binned Jensen-Shannon Divergence for perception for such scenarios to better measure the effectiveness of automatic evaluations. Third, we present visualization techniques – perception charts, to contextualize correlation measures appropriately and to show the strengths and limitations of automatic evaluation. We have open-sourced our analysis and visualization tools at https://github.com/amazon-science/BeyondCorrelation.
@@ 1 Introduction
Despite the increasing importance of automatic evaluation for generative large language models (LLMs), measuring and interpreting the effectiveness of these methods remains a challenge. Automatic evaluation methods include $N$-gram-based methods such as Rouge (Lin, 2004) and BERTScore (Zhang* et al., 2020), task-specific model-based methods such as AlignScore (Zha et al., 2023) and Prometheus (Kim et al., 2024), and LLM-as-a-Judge-based methods (LJ) where a general purpose LLM, using prompts, is instructed to perform evaluation. The use of automatic evaluation is tied to the key question – How can the efficacy of automatic evaluation methods be measured in comparison to human evaluation? A widely accepted practice is to use agreement or rank correlation metrics such as Krippendorff’s-$\alpha$ (Krippendorff, 2011), Cohen’s-$\kappa$ (Cohen, 1960; McHugh, 2012), Spearman’s $\rho$, and Kendall’s $\tau$-b (Fabbri et al., 2021; Deutsch et al., 2022; Chiang & Lee, 2023; Liu et al., 2023b), where a higher correlation between machine labels (generated by automatic evaluation methods) and human labels (gathered during human evaluation) implies a more effective automatic evaluation. Variation or uncertainty in human labels (Tversky & Kahneman, 1974) is almost impossible to avoid when humans are involved (Kahneman et al., 2021). In this paper, we highlight how comparing the overall correlation score between human labels and machine labels can obscure key discrepancies in machine labels. Therefore, we propose that measurements be stratified by label uncertainty.
@@
The impact of human label uncertainty on automatic evaluation, therefore, leads to a follow-on question – how then to effectively account for human uncertainty in automatic evaluation. Not all types of uncertainties are equal (Kahneman & Tversky, 1982), as they can range from random human errors, ambiguity in tasks, and a by-product of perception-based evaluation (Elangovan et al., 2024). Some human evaluation tasks naturally have higher variance than others. For instance, when evaluating aspects such as “How readable is the given content?”, the answer relies on human perception and is likely to vary from person to person. In contrast, evaluating a natural language inference (NLI) task, such as “Can you infer the given sentence from a given text?”, is less likely to show variations. Acceptable variation in labels, therefore, depends on the task, as variations or uncertainties in human labels can sometimes be a feature rather than a bug in human evaluation. Thus, we explore how to effectively measure and interpret reliability of automatic evaluation while taking human uncertainty in to consideration. We ask the following research questions:
@@
  • RQ1: How does uncertainty in human labels impact correlation metrics when we measure the efficacy of automatic evaluation methods? (See detailed analysis on stratifying data by human uncertainty in Sec. 3.1 )
@@
  • RQ2: How can we measure human-machine agreement that accounts for human uncertainty as a result of variation in human perception? (See Binned Jensen-Shannon divergence in Sec. 3.1 )
@@
  • RQ3: How can we visualize the underlying data to draw meaningful insights when we compare the results from automatic and human evaluations? (See Perception charts in Sec. 3.3 )
@@ 2.1 Correlation measurements and assumptions
Correlation measures have traditionally been used to measure the inter-rater agreement (IRA) of human-generated labels, where there is no ground truth available, to measure human-to-human (HH) correlation. There are 2 broad categories of correlation measures, (a) non-chance-adjusted indices and (b) chance-adjusted metrics (Xinshu Zhao & Deng, 2013). Chance-adjusted measures take into account a scenario where raters can choose the same label by chance even when they do not know the correct answer. Chance-adjusted measures, such as Krippendorff’s-$\alpha$ are generally more popular IRA measures than non-chance-adjusted metrics such as percentage agreement. The key difference between various chance-adjusted IRA measures is how chance agreement is computed (Meyer et al., 2014; Xinshu Zhao & Deng, 2013). For instance, Randolph’s-$\kappa$ (Randolph, 2005) assumes that all labels are equally likely and assigns a probability of $\nicefrac{{1}}{{k}}$ for chance agreement, where $k$ is the number of possible label choices. On the other hand, Fleiss’s $\kappa$ (Fleiss, 1971) and Krippendorff’s $\alpha$ calculate chance agreement empirically, based on the actual human labels of the study. These measurements can be useful in cases where certain labels (e.g., 3 on a Likert scale of 1-5) are more likely to be chosen when humans are uncertain. A high proportion of such labels, even if consistent, may indicate unreliability in the labeling process, particularly when the minority labels show lower agreement. Therefore, chance-adjusted correlation measures rely on key assumptions about the types of errors humans are likely to make. Furthermore, Krippendorff’s-$\alpha$, specifies that the agreement coefficient is an indicator of reliability if and only if (1) the raters duplicate the process of rating, i.e., when two independent raters use the same guidelines or instructions and use the same units, (2) raters must be treated as interchangeable, and (3) correct answers are not known (Krippendorff, 2004). Human labels and machine labels are clearly not interchangeable even with current state-of-the-art LLMs to generate machine labels, as humans provide the ground truth labels.
@@ 2.2 Differences between correlation measures and raw accuracy
The assumptions above, especially the assumption that raters must be treated as interchangeable, raise a fundamental question as to why even use correlation measures to compare machine and human performance, instead of measures such as accuracy and F1 using human labels as ground truth? Firstly, one of the main challenges with using scores like accuracy or F1 is that human labels tend to be noisy regardless of the task or the dataset, where some label choices are more acceptable than others, and measures such as F1 or accuracy do not account for variations in human labels and hence correlation measures appear to be a better measure. However, as mentioned previously, many of the common chance-adjusted correlation measures aim to account for random chance agreement in HH agreement, whereas the errors machines make tend to be more systematic. A key characteristic of random errors is that they are non-reproducible. When repeating the labeling process, the result varies, resulting in low IRA, which in-turn may point to potential problems in the labeling process or complexities in the task. In contrast, systematic errors are consistent, and IRA measures are not designed to detect them. Secondly, as a direct consequence, IRA metrics assume that agreement is good and variance is bad, which implies that there is a single correct answer – a situation that may not hold for ambiguous items or those that gauge user perception.
@@ 2.3 Types of raw data collected during evaluation
The type of the data collected, such as nominal, ordinal, interval, ratio, discrete or continuous data, dictate how such data can be aggregated and interpreted. The most common types of values collected during automatic or human evaluation are:
@@
Ordinal values: Ordinal values have an implicit order, e.g., ratings of quality (very good, good, fair, poor, very poor) (Mishra et al., 2018). The most common use of ordinal values in human evaluation are through Likert scales. Likert scales are rating systems of attitude scale to measure the degree to which a person agrees/disagrees with a given statement (Taherdoost, 2019) and commonly used to gauge user perception on qualitative aspects such as readability, coherence, and fluency (van der Lee et al., 2021). IRA measures for ordinal values should ideally account for the fact that ratings that are closer to one another (1 vs. 2) for a given item indicate a higher agreement than ratings that are further apart (1 vs. 4). Such IRA measures include Krippendorff’s-$\alpha$ for ordinal values, where disagreements are weighted by distance between the ratings. Rank correlations such as Kendall’s-$\tau$ and Spearman’s-$\rho$ are also widely used to measure item-level correlation between machines and humans for ordinal values. However, there’s a caveat on using these measures: their ability to handle a relatively large proportion of ties is widely debated, even when measuring system correlations that involve a much smaller proportion of ties (Deutsch et al., 2023). When the number of items to rank is $n$, the number of ordinal values to assign is $r$, and when $r<<n$, it results in many ties or rank collision. For instance, when assigning a Likert scale of 1-5 to 100 items, at least 20% of the items will have the same rank. While ordinal values preserve order, the objective of using ordinal values is not to rank the item in relation to other items, therefore the proportion of ties affects the interpretability and reliability of the rank correlation scores.
@@ 3 Analysis Settings and Results
Datasets and human labels: We analyze typical evaluation tasks for LLM-generated text, including (a) Likert-based qualitative evaluation (ordinal), (b) NLI-style fact or consistency checking (nominal), and (c) human preferences of several LLMs (nominal). We select datasets with multiple human annotations: 1. SummEval (Fabbri et al., 2021) is human evaluation of summary quality using four criteria – fluency, coherence, consistency, and relevance, with Likert scores assigned from 1 to 5. This dataset has 1600 samples, where each item is labeled 3 times by expert annotators. 2. SNLI (Bowman et al., 2015) is an NLI inference task dataset with 10K samples and 3 labels – entailment, neutral and contradiction. Each item is labeled 5 times by human annotators. 3. MNLI (Williams et al., 2018) is also an NLI inference task dataset with 10K samples and 3 labels – entailment, neutral and contradiction. Each item is assigned 5 labels by human annotators. It contains two splits – matched and mismatched. 4. MT-Bench (Zheng et al., 2023) is a preference dataset which compares human preferences of several model outputs on open-ended questions using multi-turn conversations. From this dataset, we only use samples that have at least 3 human ratings. 5. We also include JudgeBench (Bavaresco et al., 2024) in Appendix A.12.
Machine labels: For SummEval, we reuse the original G-Eval results (Liu et al., 2023b) which rates the quality of summaries on a scale of 1-5. In addition, we use Claude Sonnet 3 and Mistral Large v2 on SummEval, where prompts are detailed in Appendix A.2. For SNLI and MNLI, we use Llama 3.1 8B, Mistral Large v2 and Sonnet 3, reusing the prompts from Liu et al. (2023a) as detailed in Appendix A.1. For MT-Bench, we rely on the existing GPT-4 labels in the dataset (Zheng et al., 2023). We also adapt the same prompts (see Appendix A.3) for Sonnet 3 as LJ. In order to compute MM correlation, we set the sampling temperature to 1.0 to maximize the variation across runs, and set top_p = 0.8. We sample 20 machine labels per example. NOTE: We have predominately reused the prompts from existing works of using GPT-4 as LJ, we have NOT attempted to optimize the prompts per LLM except for minor format changes. Since models are known to be highly sensitive to prompts, our experiments are not meant to rank the underlying LJ, but to study the impact of human uncertainty regardless of the LJ.
@@
To study the impact of human uncertainty, we stratify samples and then compare HH and Human-to-machine (HM) correlation measurements for each group. We primarily stratify by percentage agreement, which is the proportion of labels that matches the majority (for nominal data) or the median label (for ordinal data) among all human labels assigned to an item. For example, if 3 out of 5 people assign the same label, the percentage agreement is 60%. We also stratify samples by the number of unique human labels to ensure that our findings are consistent regardless of the stratification method.
In Table 3.1, column $\text{H}^{w}\text{M}^{w}$ compares Human majority ($\text{H}^{w}$) with machine majority ($\text{M}^{w}$). Superficially, HM correlation seems to improve with increasing human certainty. On closer inspection, another trend emerges where in any stratified group, when uncertain samples increases (as measured by low HH correlation), the HM correlation seems to outperform HH correlation as highlighted. However, as the proportion of samples with noisy labels decreases and HH correlation approaches near-perfect agreement, HH correlation is much higher than HM as highlighted. Thus, the illusion that machines (specifically LJs) approximate majority human behavior is largely due to the presence of noisy labels in the data. If LJs were truly approximating the human majority, the $\Delta$ would approach 0 under perfect HH correlation – which is clearly not the case. We also replicated this behavior on SummEval dataset which relies on ordinal labels in Table 2, using correlation metrics such as Krippendorff’s-$\alpha$ for ordinal data, Kendall’s-$\tau$ and Spearman’s-$\rho$. The same patterns are observed on preference data in Table 3. We further randomly sample subsets from NLI datasets (sampling detailed in Appendix A.5) and compare correlation scores across subsets with varying levels of human label uncertainty. We repeat this for three different LLMs, and consistently observe the same findings, performance gaps become more apparent as human label consistency increases, as shown in Fig. 1. Similar findings for the SummEval dataset are shown in Appendix Fig. 5.
We explore the phenomenon using a random labeler on synthetically labeled datasets (code provided in Appendix A.11). This dataset simulates both binary classification and ordinal classification with labels {1, 2, 3}. We first sample 2 synthetic human labels to each item. Then another random labeler assigns labels randomly. We examine the correlation between the random labels and the synthetic human labels across varying levels of certainty. Notably, the random labeler may appear to have higher correlation when the proportion of uncertain samples is higher. However, as the proportion of consistently labeled samples increases, the random labels’ poor performance becomes evident, as shown in Fig. 2. Intuitively, if 2 human labels disagree, a 3rd random label cannot perform worse in terms of agreement – they can either disagree with both human labels (in which case it is no better than the 2 humans) or agree with one of the human label. This example further illustrates why stratification by proportion of uncertainty is crucial to show weaknesses in any automatic labeler.
@@
Table 1: Compare HM and HH correlation measurements on NLI datasets using Sonnet to generate machine labels. The overall correlation scores (All) indicate that the HM correlation is similar to HH. However, when stratifying by percentage agreement (PA) or human unique labels, the difference ($\Delta=\text{HH}-\text{H}^{w}\text{M}^{w}$) is highest when the all human labels are the same (PA=1), and $\text{H}^{w}\text{M}^{w}$ outperforms HH when certainty is the lowest. $\text{H}^{w}$ and $\text{M}^{w}$ are human and machine majority label.
@@
Figure 1: Krippendorff’s $\alpha$ (K-$\alpha$) on subsets of MNLI-mismatched with varying uncertainty, using 3 LLMs – Sonnet, Mistral, and Llama. As proportion of samples with consistent human labels increases (towards the right of X-axis), HH correlations are significantly higher than $\text{H}^{w}\text{M}^{w}$. $\text{H}^{w}\text{H}^{\prime}$ represents the correlation between the human majority and another human annotator ($\text{H}^{\prime}$). $\text{H}^{w}\text{R}^{w}$ compares the human majority with a random machine evaluator, serving as baseline performance.
Figure 2: Simulating the impact of uncertainty by comparing with an automatic random labeler (R). When the proportion of samples with uncertainty is higher, even a random labeler can appear to have better correlation with a majority ($\text{H}^{w}$) or median ($\overline{\text{H}}$) human label. As the proportion of samples with consistency increases, the weakness of the random labeler become evident.
@@
Table 2: Results on SummEval dataset for criteria – Coherence and Relevance, $\overline{\text{H}}$ is the human median rating and $\overline{\text{M}}$ is machine median. $\Delta=(\text{HH}-\overline{\text{H}}\overline{\text{M}}$). Impact of human uncertainty: When the human label uncertainty is high, $\overline{\text{H}}\overline{\text{M}}$ can seem better than HH correlation and vice versa under low uncertainty. Does the correlation score ($\rho$ of 0.56 , $\alpha$ 0.49) imply the LJ is good enough to replace human evaluation? Spearman’s-$\rho$ median $(\overline{\text{H}}\overline{\text{M}}$) vs. mean ($\text{H}^{\mu}\text{M}^{\mu}$): Median vs. mean results in an 11 points boost 0.49 $\rightarrow$ 0.60. Differences in correlation metrics: Krippendorff’s-$\alpha$ of 0.25 (reject results), while Spearman’s-$\rho$ shows 0.53 (moderate correlation). LJ ranking and absolute correlation numbers change depending on reference human under high uncertainty, Mistral best performance 0.42 replaced by Sonnet 0.22 when reference human is $\text{H}^{\prime}$ (another human annotator). Full table in Appendix Table 7.
@@
Table 3: Stratified performance on preference data with at least 3 human ratings from MT-Bench for model pairs (MP). Impact of human label uncertainty: When the HH label uncertainty is high, $\text{H}^{w}\text{M}^{w}$ can seem better than HH correlation and vice versa. IRA vs. Win-rate: While the difference in win-rate between model and human ($\Delta$) can be smaller in one evaluator model (e.g., Sonnet 0.08) compared to another (e.g., GPT 0.21), the $\text{H}^{w}\text{M}^{w}$ IRA can tell an opposite story, as in the example above, GPT has better correlation with human majority across all correlation metrics except $JS_{b}$. LJ ranking and absolute correlation numbers change depending on reference human $\text{H}^{\prime}$, mimicking single annotator, under high uncertainty compared to using consensus human majority ($\text{H}^{w}$).
3.2 RQ2: How can we measure human-machine agreement that accounts for human uncertainty as a result of variation in human perception?
@@
where for ordinal values, bin $\mathbf{B}_{r}\coloneqq\{i\in\{1,..n\}|\text{median}(H_{i})=r\}$ contains the set of samples with a median human rating $r$. For nominal perception-based preference labels, bin $\mathbf{B}_{r}\coloneqq\{i\in\{1,..n\}|\text{mostfrequent}(H_{i})=r\}$ contains the set of samples with the most frequent human rating $r$.